OpenStudy (anonymous):

Given the function f(x) = log4(x + 8), find the value of f−1(2). a. f−1(2) = 4 b. f−1(2) = 8 c. f−1(2) = 10 d. f−1(2) = 17

OpenStudy (anonymous):

hi are you still on? i can help

OpenStudy (anonymous):

we need to find the inverse. do you know how to do that?

OpenStudy (anonymous):

don't i replace x with y and y with x?

OpenStudy (anonymous):

f(x) = log4(x + 8) y = log4(x + 8) x = log4(y + 8) Now solve for y from there. x = log4(y + 8) \[f-1(x) = -8 + 4^x\]

OpenStudy (anonymous):

now just plug in 2, what do you get?

OpenStudy (anonymous):

-8 + 4^2 4^2 = 16 -8+16 = 8 your answer will be B.

OpenStudy (anonymous):

do you need anymore help?

OpenStudy (anonymous):

if so close this question and post another one and invite me. please medal me though

OpenStudy (anonymous):

oh okay, thank you so so much! and yes, i do. is it okay if i post it separately? so incase any one else needs help, they can find it easier.

OpenStudy (anonymous):

yes awesome